Leianna:

A cone with a radius of 6 cm has a volume of 120π cm3. What is the height of the cone? (The volume of a cone is given by the formula V = 1 3 πr2h.)

ramen:

try working backwards to find the height

TheSmartOne:

Hey there! Plug in r = 6 into the equation that you're given for the volume of the cone and v = 120\(\pi\) and solve for h \(\ V = \frac{1}{3}\pi r^2 h\)

Leianna:

okay thank you!

TheSmartOne:

Let us know what you get and we can check for you!! Also welcome to QuestionCove!! We're glad to have you here

ramen:

120pi=1/3pi(6)^2h, then solve for h

Leianna:

I really appreciate you guys

ramen:

no problem

Leianna:

I got an 11.2

TheSmartOne:

Can you show us your work or what you put into the calculator?

TheSmartOne:

\(\ 120\pi = \frac{1}{3}\pi (6)^2 h\)

Leianna:

actually I got 10.47

TheSmartOne:

All the stuff on the left hand side are being multiplied so to get rid of them you'd have to divide So like you'd divide pi on both sides and that would cancel that out and so on

Leianna:

I did 120/6^2(pi)

TheSmartOne:

so we have 120 pi! And you forgot the 1/3 So let's do it one step at a time \(\color{#0cbb34}{\text{Originally Posted by}}\) @TheSmartOne \(\ 120\pi = \frac{1}{3}\pi (6)^2 h\) \(\color{#0cbb34}{\text{End of Quote}}\) What happens when you divide pi on both sides? Will it cancel out or not?

Leianna:

Yes

TheSmartOne:

Perfect! So now we have \(\ 120= \frac{1}{3} (6)^2 h\)

TheSmartOne:

What is 6^2 equal to?

Leianna:

36

TheSmartOne:

Good good Now what is 1/3 * 36

Leianna:

12

TheSmartOne:

Good So now we have \(\ 120= 12 h\)

TheSmartOne:

Can you solve for h now?

Leianna:

120/12

TheSmartOne:

Yes and simplify that now

Leianna:

How do I do that?

TheSmartOne:

Well how would you make it smaller Like what is 120/12 You should be able to do it in your head if you can see 120 is 12*10 Or you could use a calculator

Leianna:

it is 10

TheSmartOne:

Simplify basically means to bring into smaller numbers So like if you had 24/4 we could "simplify" it to (12*2) / (2*2) And we can cross out a 2 from the numerator and denominator And we get 12/2 But we can simplify it more!! (6*2) / (2) And that ultimately is equal to 6

TheSmartOne:

Yes, 10 is correct!!

Leianna:

thank you so so so much

ramen:

yep good job thats what i got too

TheSmartOne:

Of course!! Hopefully that made sense! Just remember to next time do it step by step so that way you don't make any silly mistakes

ramen:

oh yea when you work backwards remember sadmep

Leianna:

I will remember that

Leianna:

Seriously guys thank you so very much

TheSmartOne:

Of course! That's why we're here!

ramen:

no problem